OpenStudy (anonymous):

Solve for x x^2+37=2x Solutions are?

OpenStudy (anonymous):

x^2 - 2x + 37 = 0 a = 1 b = -2 c = 37 D = b-4ac x = (-b +- sqrt (D)) / 2a

OpenStudy (anonymous):

are you sure it's okay? suzi

OpenStudy (anonymous):

x = 1 - 6i x = 1 + 6i

OpenStudy (anonymous):

that was my answer!

OpenStudy (anonymous):

D = b^2 -4ac

OpenStudy (anonymous):

:D